# bywithappendix

### v1

@@ 1. Introduction
In fixed-income markets, the basic financial instruments are securities that are based on interest rate payments. A common example among these are government bonds, which typically entail the holder to a fixed coupon, which is a regular payment, and an additional nominal at the end of the contract. The two most principle approaches to model all of the government bonds at once are the Heath-Jarrow-Morton (HJM) approach [HJM92] and the LIBOR market model [BGM97]. In the HJM approach an underlying infinite-dimensional process $(f(t,T))_{0\leq t\leq T}$ is modelled, where $f(t,T)$ is called the forward rate at time $t$ and time of maturity $T$. In this framework, the zero-coupon bond price at time $t$ with contract maturity date $T$ is defined via
@@
It is convenient to use (with slight abuse of notation) the parametrisation $f_{t}(x):=f(t,t+x)$ due to [Mus93] which expresses the forward rate in terms of time and time-to-maturity instead of time and maturity. One can consider the forward rate as a curve-valued object and arrive at the Heath-Jarrow-Morton-Musiela (HJMM) formulation. The original HJM-drift condition which ensures absence of arbitrage has been reformulated in [Mus93] into the time-to-maturity formulation.
@@
In order to make these models computable, one possibility is to set up a finite factor model $(Z_{t})_{t\geq 0}$ with $f_{t}=g(Z_{t},\cdot)$ for some deterministic function $g$. A more principled approach to this is discussed in [Bjö04] where, instead of a finite-parameter set of curves, a manifold $\mathcal{M}$ of curves is used. [Bjö04, p. 17] asks three questions which are all about the relation of a forward rate curve model $(f_{t})_{t\geq 0}$ and some manifold of curves $\mathcal{M}$, as for instance: Under which condition does a given forward rate curve model stay on a prescribed manifold of curves?
@@
From a more general standpoint, manifolds which remain invariant under (stochastic) dynamical systems are an important object of study as a solution space of (stochastic) partial differential equations ((S)PDEs). In mathematical finance those invariant manifolds arise naturally when considering finite-dimensional realisations of (potentially infinite-dimensional) consistent term structure models (see, e.g. [BC99, BL00, BLS02, BS01, FT01, Tap05, Tap15, Tap10]). There are several motivations for considering finite-dimensional realisations. For instance, finite dimensional models are easier simulated in a computer and if there is a finitely parametrised family of curves it is easy to statistically decide for a curve from data than in some infinite dimensional case.
@@
We are interested in the question of when, for a given manifold of curves $\mathcal{M}$, a model for the forward rate curve can be constructed. This problem was analysed in [Fil00], where equivalent conditions have been derived. In particular, any diffusion model for $f$ which is compatible with a given manifold $\mathcal{M}$ must have at least a diffusion coefficient which is tangential to $\mathcal{M}$. Additional results, as in e.g. [FTT14] further characterise invariant finite-dimensional manifolds for SPDEs describing the evolution of the forward curve $f$. Examples of finite-dimensional manifolds which are invariant under consistent HJMM dynamics are known. Most notably, affine term structures which are finite dimensional affine spaces have been studied in e.g. [FT03, FT04] where the authors show that the only HJMM-type models which admit a finite-dimensional realisation for essentially any initial curve $f_{0}$ are affine in the direction of the diffusion coefficient. In contrast to our problem specification, the authors construct invariant (finite-dimensional) manifolds given a diffusion coefficient, which turn out to be necessarily affine.
@@
The problem of calibrating financial models is well known in mathematical finance. In particular, calibration of term structure models has to be done in a careful way to preserve no-arbitrage conditions specified by the respective model. Recalibrating models frequently may lead to violations thereof. In [RT17] the authors consider discrete-time recalibration schemes which generate arbitrage-free models for any time via so-called forward characteristics. Applications for term structure models of bonds and volatility curves are considered. To calibrate the term structure of interest rates, often times parametric families of curves are calibrated to properly interpolate observed prices. It is well-known (see, e.g. [Fil01]) that these families have to be chosen carefully, as otherwise dynamic consistency fails to be maintained. More recent approaches have been studied, as in e.g. [WJ24], where the authors attempt to employ a smoothing procedure for cubic splines to fit yield curves. To maintain dynamic consistency, the authors add higher-degree polynomials to the linear span of functions. In [LMS24], the authors use deep-learning methods to obtain a low-dimensional representation of the term structures via autoencoders. These, however, do not necessarily stay on the invariant manifold generated by the consistent model and instead remains within a vicinity dependent on the volatility.
@@ 2.1. Notation
    $\mathcal{\operatorname{S}}_{h}:\mathcal{H}\rightarrow\mathcal{H},f\mapsto(x\mapsto f(x+h))$ defines a $c_{0}$-semigroup $\mathcal{(}\operatorname{S}_{h})_{h\geq 0}$ on $\mathcal{H}$ whose generator will be denoted by $\partial_{x}$, cf. [EK09, pp. 6–8].
We denote the scalar product on $\mathcal{H}$ by $\langle h_{1},h_{2}\rangle_{\mathcal{H}}$ for $h_{1},h_{2}\in\mathcal{H}$. We will make use of a Hilbert space $\mathcal{H}$ satisfying Assumptions LABEL:H1 and LABEL:H2 throughout the paper without explicitly referencing them. An example of a Hilbert space with those properties is the forward curve space $\mathcal{H}_{w}$ (see [Fil01]) defined as follows:
@@ 2.2. Setting
We are interested in characterising the finite dimensional manifolds $\mathcal{M}\subseteq\mathcal{H}$ which are invariant under the HJMM-equation relative to any diffusion coefficient which is tangential to the manifold. The motivation of the problem comes from a statistical consideration. If one has decided for a set of curves $\mathcal{M}$ and estimates the diffusion coefficient of a process running on $\mathcal{M}$, then we like to be sure to obtain a model which is free of arbitrage in the sense of NAFLVR (see [CKT16]). Indeed, it is desirable for a calibrated model to remain within the admissible space of functions no matter the values of the estimated parameters. We proceed by making mathematically clear what we mean and summarise our main finding in Theorem 3.3 below. We will assume that the manifold is embedded into a Hilbert space of functions which is also used to formalise the HJMM-equation.
@@
The map $\chi$ is called a local parametrisation of $\mathcal{M}$. We shall denote by $T_{m}$ the tangent space of $\mathcal{M}$ at $m\in\mathcal{M}$, that is $T_{m}:=\{D\chi(y)v:v\in\mathbb{R}^{d}\}$ for $y\in U$ so that $\chi(y)=m$. For the exact definition we refer to [FTT14a, Section 3].
@@
    $\Sigma$ is measurable and locally bounded in the sense of [Tap12, Definition 3.5].
@@
in the sense of [DZ92, Section 6.1], where $\beta$ is as in Definition 2.4.
Following the arguments in [Zab00], the study of invariant manifolds is interesting for the sake of model calibration. Indeed, for curve fitting for interest rate models, it is desirable for the stochastic models to stay within a prescribed manifold while still staying consistent with the NAFLVR condition. This is made precise in the following definitions.
@@
Let $\mathcal{M}$ be a $C^{2}$-manifold, $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ be of class $C^{1}$ and $\beta:\mathcal{H}\rightarrow\mathcal{H}$ locally Lipschitz and locally bounded in the sense of [Fil01, Definition 2.4.2, Definition 2.4.3], respectively. [Fil00, Theorem 3], [FT03] have shown that the following statements are equivalent:
@@
Moreover, [Fil00, Theorem 2] yields that if $\mathcal{M}$ is closed in $\mathcal{H}$, then invariance is ensured for $\tau=\infty$. [FTT14] extend this to a setting with jump diffusions with the additional requirement of $\mathcal{M}$ being of class $C^{3}$.
@@ 3. Main results
The difference between our setting and the setting in [FT03, FT04] is as follows. In [FT04, Theorem 5.3] they answer the question: Given an admissible volatility structure, which HJMM-models permit a finite-dimensional realisation for essentially any initial curve $f_{0}$. Indeed, any invariant, finite-dimensional manifold constructed for such an initial curve $f_{0}$ turns out to be affine in the direction of the volatility. Our result answers a different question, namely: Which manifolds $\mathcal{M}$ are invariant for any tangential diffusion coefficient $\Sigma$ and any initial curve $f_{0}\in\mathcal{M}$? Here, the construction already assumes that for any $f_{0}\in\mathcal{M}$, there is a solution for the HJMM-equation with diffusion $\Sigma$ which is $\mathcal{M}$-valued, and our result reveals the structure of these manifolds. It turns out that these are of a linear-rational nature, cf. Theorem 3.3.
@@
The Hilbert space $\mathcal{H}_{w}$ given in Definition 2.1 with $\mathcal{H}_{w}^{0}\subseteq\mathcal{H}_{w}$ given by [Fil01, Equation 5.12] satisfies that $f_{h}$ as given in the previous theorem is in $\mathcal{H}_{w}$ for any $h\in\mathcal{H}_{w}^{0}$ as shown in [Fil01, Lemma 5.2.1].
@@ 4. Proof of the main statement
Since $\mathcal{H}$ fulfills Assumption LABEL:H2, $\operatorname{S}$ is strongly continuous. This implies that the maps $\phi_{h,h^{\prime}}$ are continuous for all $h,h^{\prime}i\in\mathcal{H}$. This implies that the adjoint semigroup $(\operatorname{S}_{x}^{*})_{x\geq 0}$ is weakly continuous. By [EN00, Theorem 5.8], $\operatorname{S}^{*}$ is strongly continuous (we note that more generally this holds on any reflexive Banach space) and accordingly, the map $x\mapsto k_{x}$ is continuous and therefore (locally) integrable. Thus, $\beta$ is well-defined. Since $\Sigma$ is Lipschitz and with bounded support, $\beta$ is Lipschitz. [DZ92, Theorem 7.2] yields existence of the process $(f_{t})_{t\geq 0}$. ∎
@@
Let $h\in\mathcal{M}$ and let $U\subseteq\mathcal{H}$ be an open neighbourhood of $h$. Choose $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ to be of class $C^{1}$ with support in $U$ and $\Sigma(m)\in T_{m}$ for any $m\in U\cap\mathcal{M}$. Since $\mathcal{M}$ is totally invariant we find that $\mathcal{M}$ is invariant for the HJMM-equation with diffusion coefficient $\Sigma$. Hence, [Fil00, Theorem 3] yields that we have
@@
The claim now follows by the same arguments as [KX23, Lemma 2.9]. ∎
@@
The proof of the above Lemma is an adaptation of the proof in [KX23, Proposition 2.14] to our case.
@@
    Theorem 3.3 yields the asserted manifold $\mathcal{M}$ for some $C^{2}$ functions $c:\mathbb{R}_{+}\rightarrow\mathbb{R}$ and $u:\mathbb{R}_{+}\rightarrow\mathbb{R}$. Using the results of [Fil23, Proposition 3.2] we obtain a matrix $M\in\mathbb{R}^{(d+1)\times(d+1)}$ such that $c$ and $u$ assume the asserted form.
@@
    Since $c$ and $u$ are real analytic functions, this easily implies $V\subseteq\text{dom}(\partial_{x})$, that is 2.9.ii)a. Now, consider the pair $(\beta^{h},\Sigma^{h})\in\Lambda^{h}$. Since $\operatorname{J}_{\cdot}$ is a linear operator, $\Sigma^{h}(h)\in L(\mathbb{R}^{d},T_{h})$ for any $h\in V$, which is directly 2.9.ii)b of. To verify the third condition, we observe that $V$ is an affine manifold. Due to [Fil00, Theorem 4], this reduces Condition 2.9.ii)c to $\partial_{x}h+\beta^{h}(h)\in T_{h}$ for any $h\in T_{h}$. To show that this is true, define $r:=\delta_{0}(\Psi^{-1}h)$. Due to Lemma 4.1, we find that $r=h^{\prime}(0)$. Using the particular form of the functions $c$ and $u$, we therefore have $r=\langle-Me_{1},(1,z)\rangle$ for $z\in U$. We may now compute
@@ Appendix A Technical tools
as $n\rightarrow\infty$ because [EK09, Ch. 1, Proposition 1.2] yields that $\sup_{x\in[0,x_{1}]}\|\operatorname{S}_{x}\|<\infty$. ∎

### v2

@@ 1. Introduction
In fixed-income markets, the basic financial instruments are securities that are based on interest rate payments. A common example among these are government bonds, which typically entail the holder to a fixed coupon, which is a regular payment, and an additional nominal at the end of the contract. The two most principle approaches to model all of the government bonds at once are the Heath-Jarrow-Morton (HJM) approach [hjm] and the LIBOR market model [brace_al_LIBOR]. In the HJM approach an underlying infinite-dimensional process $(f(t,T))_{0\leq t\leq T}$ is modelled, where $f(t,T)$ is called the forward rate at time $t$ and time of maturity $T$. In this framework, the zero-coupon bond price at time $t$ with contract maturity date $T$ is defined via
@@
It is convenient to use (with slight abuse of notation) the parametrisation $f_{t}(x):=f(t,t+x)$ due to [musiela] which expresses the forward rate in terms of time and time-to-maturity instead of time and maturity. One can consider the forward rate as a curve-valued object and arrive at the Heath-Jarrow-Morton-Musiela (HJMM) formulation. The original HJM-drift condition which ensures absence of arbitrage has been reformulated in [musiela] into the time-to-maturity formulation.
@@
In order to make these models computable, one possibility is to set up a finite factor model $(Z_{t})_{t\geq 0}$ with $f_{t}=g(Z_{t},\cdot)$ for some deterministic function $g$. A more principled approach to this is discussed in [bjoerk] where, instead of a finite-parameter set of curves, a manifold $\mathcal{M}$ of curves is used. [bjoerk, p. 17] asks three questions which are all about the relation of a forward rate curve model $(f_{t})_{t\geq 0}$ and some manifold of curves $\mathcal{M}$, as for instance: Under which condition does a given forward rate curve model stay on a prescribed manifold of curves?
@@
From a more general standpoint, manifolds which remain invariant under (stochastic) dynamical systems are an important object of study as a solution space of (stochastic) partial differential equations ((S)PDEs). In mathematical finance those invariant manifolds arise naturally when considering finite-dimensional realisations of (potentially infinite-dimensional) consistent term structure models (see, e.g. [bjoerk1, bjoerk2, bjoerk3, bjoerk_svensson, filip, Tappe, tappe2, tappe4]). There are several motivations for considering finite-dimensional realisations. For instance, finite dimensional models are easier simulated in a computer and if there is a finitely parametrised family of curves it is easy to statistically decide for a curve from data than in some infinite dimensional case.
@@
We are interested in the question of when, for a given manifold of curves $\mathcal{M}$, a model for the forward rate curve can be constructed. This problem was analysed in [filipovic_invariant], where equivalent conditions have been derived. In particular, any diffusion model for $f$ which is compatible with a given manifold $\mathcal{M}$ must have at least a diffusion coefficient which is tangential to $\mathcal{M}$. Additional results, as in e.g. [tappe_invariant] further characterise invariant finite-dimensional manifolds for SPDEs describing the evolution of the forward curve $f$. Examples of finite-dimensional manifolds which are invariant under consistent HJMM dynamics are known. Most notably, affine term structures which are finite dimensional affine spaces have been studied in e.g. [teichmann1, teich] where the authors show that the only HJMM-type models which admit a finite-dimensional realisation for essentially any initial curve $f_{0}$ are affine in the direction of the diffusion coefficient. In contrast to our problem specification, the authors construct invariant (finite-dimensional) manifolds given a diffusion coefficient, which turn out to be necessarily affine.
@@
The problem of calibrating financial models is well known in mathematical finance. In particular, calibration of term structure models has to be done in a careful way to preserve no-arbitrage conditions specified by the respective model. Recalibrating models frequently may lead to violations thereof. In [richter_teichmann] the authors consider discrete-time recalibration schemes which generate arbitrage-free models for any time via so-called forward characteristics. Applications for term structure models of bonds and volatility curves are considered. To calibrate the term structure of interest rates, often times parametric families of curves are calibrated to properly interpolate observed prices. It is well-known (see, e.g. [filipovic_consistency]) that these families have to be chosen carefully, as otherwise dynamic consistency fails to be maintained. More recent approaches have been studied, as in e.g. [jarrow_wu], where the authors attempt to employ a smoothing procedure for cubic splines to fit yield curves. To maintain dynamic consistency, the authors add higher-degree polynomials to the linear span of functions. In [autoencoder], the authors use deep-learning methods to obtain a low-dimensional representation of the term structures via autoencoders. These, however, do not necessarily stay on the invariant manifold generated by the consistent model and instead remains within a vicinity dependent on the volatility.
@@ 2.1. Notation
    $\mathcal{\operatorname{S}}_{h}:\mathcal{H}\rightarrow\mathcal{H},f\mapsto(x\mapsto f(x+h))$ defines a $c_{0}$-semigroup $\mathcal{(}\operatorname{S}_{h})_{h\geq 0}$ on $\mathcal{H}$ whose generator will be denoted by $\partial_{x}$, cf. [EK, pp. 6–8].
We denote the scalar product on $\mathcal{H}$ by $\langle h_{1},h_{2}\rangle_{\mathcal{H}}$ for $h_{1},h_{2}\in\mathcal{H}$. We will make use of a Hilbert space $\mathcal{H}$ satisfying Assumptions LABEL:H1 and LABEL:H2 throughout the paper without explicitly referencing them. An example of a Hilbert space with those properties is the forward curve space $\mathcal{H}_{w}$ (see [filipovic_consistency]) defined as follows:
@@ 2.2. Setting
We are interested in characterising the finite dimensional manifolds $\mathcal{M}\subseteq\mathcal{H}$ which are invariant under the HJMM-equation relative to any diffusion coefficient which is tangential to the manifold. The motivation of the problem comes from a statistical consideration. If one has decided for a set of curves $\mathcal{M}$ and estimates the diffusion coefficient of a process running on $\mathcal{M}$, then we like to be sure to obtain a model which is free of arbitrage in the sense of NAFLVR (see [cuchiero_naflvr]). Indeed, it is desirable for a calibrated model to remain within the admissible space of functions no matter the values of the estimated parameters. We proceed by making mathematically clear what we mean and summarise our main finding in Theorem 3.3 below. We will assume that the manifold is embedded into a Hilbert space of functions which is also used to formalise the HJMM-equation.
@@
The map $\chi$ is called a local parametrisation of $\mathcal{M}$. We shall denote by $T_{m}$ the tangent space of $\mathcal{M}$ at $m\in\mathcal{M}$, that is $T_{m}:=\{D\chi(y)v:v\in\mathbb{R}^{d}\}$ for $y\in U$ so that $\chi(y)=m$. For the exact definition we refer to [tappe_invariant_appendix, Section 3].
@@
    $\Sigma$ is measurable and locally bounded in the sense of [tappe_existence, Definition 3.5].
@@
in the sense of [zabczyk, Section 6.1], where $\beta$ is as in Definition 2.4.
Following the arguments in [zabczyk2], the study of invariant manifolds is interesting for the sake of model calibration. Indeed, for curve fitting for interest rate models, it is desirable for the stochastic models to stay within a prescribed manifold while still staying consistent with the NAFLVR condition. This is made precise in the following definitions.
@@
Let $\mathcal{M}$ be a $C^{2}$-manifold, $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ be of class $C^{1}$ and $\beta:\mathcal{H}\rightarrow\mathcal{H}$ locally Lipschitz and locally bounded in the sense of [filipovic_consistency, Definition 2.4.2, Definition 2.4.3], respectively. [filipovic_invariant, Theorem 3], [teichmann1] have shown that the following statements are equivalent:
@@
Moreover, [filipovic_invariant, Theorem 2] yields that if $\mathcal{M}$ is closed in $\mathcal{H}$, then invariance is ensured for $\tau=\infty$. [tappe_invariant] extend this to a setting with jump diffusions with the additional requirement of $\mathcal{M}$ being of class $C^{3}$.
@@ 3. Main results
The difference between our setting and the setting in [teichmann1, teich] is as follows. In [teich, Theorem 5.3] they answer the question: Given an admissible volatility structure, which HJMM-models permit a finite-dimensional realisation for essentially any initial curve $f_{0}$. Indeed, any invariant, finite-dimensional manifold constructed for such an initial curve $f_{0}$ turns out to be affine in the direction of the volatility. Our result answers a different question, namely: Which manifolds $\mathcal{M}$ are invariant for any tangential diffusion coefficient $\Sigma$ and any initial curve $f_{0}\in\mathcal{M}$? Here, the construction already assumes that for any $f_{0}\in\mathcal{M}$, there is a solution for the HJMM-equation with diffusion $\Sigma$ which is $\mathcal{M}$-valued, and our result reveals the structure of these manifolds. It turns out that these are of a linear-rational nature, cf. Theorem 3.3.
@@
The Hilbert space $\mathcal{H}_{w}$ given in Definition 2.1 with $\mathcal{H}_{w}^{0}\subseteq\mathcal{H}_{w}$ given by [filipovic_consistency, Equation 5.12] satisfies that $f_{h}$ as given in the previous theorem is in $\mathcal{H}_{w}$ for any $h\in\mathcal{H}_{w}^{0}$ as shown in [filipovic_consistency, Lemma 5.2.1].
@@ 4. Proof of the main statement
Since $\mathcal{H}$ fulfills Assumption LABEL:H2, $\operatorname{S}$ is strongly continuous. This implies that the maps $\phi_{h,h^{\prime}}$ are continuous for all $h,h^{\prime}i\in\mathcal{H}$. This implies that the adjoint semigroup $(\operatorname{S}_{x}^{*})_{x\geq 0}$ is weakly continuous. By [engel, Theorem 5.8], $\operatorname{S}^{*}$ is strongly continuous (we note that more generally this holds on any reflexive Banach space) and accordingly, the map $x\mapsto k_{x}$ is continuous and therefore (locally) integrable. Thus, $\beta$ is well-defined. Since $\Sigma$ is Lipschitz and with bounded support, $\beta$ is Lipschitz. [zabczyk, Theorem 7.2] yields existence of the process $(f_{t})_{t\geq 0}$. ∎
@@
Let $h\in\mathcal{M}$ and let $U\subseteq\mathcal{H}$ be an open neighbourhood of $h$. Choose $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ to be of class $C^{1}$ with support in $U$ and $\Sigma(m)\in T_{m}$ for any $m\in U\cap\mathcal{M}$. Since $\mathcal{M}$ is totally invariant we find that $\mathcal{M}$ is invariant for the HJMM-equation with diffusion coefficient $\Sigma$. Hence, [filipovic_invariant, Theorem 3] yields that we have
@@
The claim now follows by the same arguments as [shijie, Lemma 2.9]. ∎
@@
The proof of the above Lemma is an adaptation of the proof in [shijie, Proposition 2.14] to our case.
@@
    Theorem 3.3 yields the asserted manifold $\mathcal{M}$ for some $C^{2}$ functions $c:\mathbb{R}_{+}\rightarrow\mathbb{R}$ and $u:\mathbb{R}_{+}\rightarrow\mathbb{R}$. Using the results of [filipovic_discount, Proposition 3.2] we obtain a matrix $M\in\mathbb{R}^{(d+1)\times(d+1)}$ such that $c$ and $u$ assume the asserted form.
@@
    Since $c$ and $u$ are real analytic functions, this easily implies $V\subseteq\text{dom}(\partial_{x})$, that is 2.9.ii)a. Now, consider the pair $(\beta^{h},\Sigma^{h})\in\Lambda^{h}$. Since $\operatorname{J}_{\cdot}$ is a linear operator, $\Sigma^{h}(h)\in L(\mathbb{R}^{d},T_{h})$ for any $h\in V$, which is directly 2.9.ii)b of. To verify the third condition, we observe that $V$ is an affine manifold. Due to [filipovic_invariant, Theorem 4], this reduces Condition 2.9.ii)c to $\partial_{x}h+\beta^{h}(h)\in T_{h}$ for any $h\in T_{h}$. To show that this is true, define $r:=\delta_{0}(\Psi^{-1}h)$. Due to Lemma 4.1, we find that $r=h^{\prime}(0)$. Using the particular form of the functions $c$ and $u$, we therefore have $r=\langle-Me_{1},(1,z)\rangle$ for $z\in U$. We may now compute
@@ Appendix A Technical tools
as $n\rightarrow\infty$ because [EK, Ch. 1, Proposition 1.2] yields that $\sup_{x\in[0,x_{1}]}\|\operatorname{S}_{x}\|<\infty$. ∎